# Green Screen Augmentation Enables Scene Generalisation in Robotic Manipulation

### v1

@@ 1 Introduction
In light of these complications, we opt for a simpler yet effective alternative: green screens. The film industry has utilised green screens extensively [14, 15, 16, 17, 18], enabling the addition of virtual backgrounds to live footage. Inspired by these applications, we apply green screen technology to robotics, allowing robots to perform tasks in unfamiliar scenes not part of the training demonstration data.
@@ 2 Related Work
Green screen in machine learning and robotics. Green screen has been traditionally used for film and video production [14, 15, 16, 17, 18]. In recent years, its application in machine learning has increased. Smirnov et al. [18] applied machine learning to improve the quality of chroma keying. Xu et al. [32], Sengupta et al. [33], Lin et al. [34, 35] explored machine learning techniques to replace green screens, enabling natural image matting without them. Schülein et al. [36] used green screens and chroma keying to replace backgrounds with clinical scenes to create synthetic data for medical clothing detection. In robotics, the use of green screens remains limited. Coates and Ng [37] employed it to develop a multi-camera object detector with synthetic data from chroma-keyed backgrounds.
@@ 3.3 Training Robot Learning Policies
GreenAug can be applied to RGB-based robot learning methods. Similar to standard augmentation methods, images can be transformed with GreenAug and fed into policy networks during training, or they can be preprocessed offline and then used for training. Offline preprocessing is more common due to the longer computation time of some GreenAug variants. However, in online settings such as reinforcement learning, online transformations are also effective. GreenAug-Rand and GreenAug-Gen allow each raw frame from the training demonstrations to be augmented with different textures, significantly increasing the amount of preprocessed data. In contrast, GreenAug-Mask only masks the background and provides a single solution. To ensure a fair comparison, we keep the number of preprocessed frames equal to the number of raw frames for all methods.
@@ 4.2 Setup
Training. We trained all baselines and our methods on both sets of data, except for GreenAug, which was excluded from Scene 2 as it relies on the green screen. Each data set corresponds to a separate policy. ACT is used as the control policy for our main experiments.
@@ 4.4 Ablation and Further Studies
Applying GreenAug to a different robot with reinforcement learning. We investigated whether GreenAug can be applied to a different robot embodiment and learning method, beyond the Franka Panda and ACT. We set up a similar “take lid off saucepan” task on a UR5. We used a continuous demo-driven DQN variant [44, 45, 46, 47] with actions discretised into bins. The robot was provided with 24 demonstrations and was given a sparse reward of 0 for failure and 1 for success. We trained the robot online with 20 minutes of exploration on a green screen background and evaluated two policies, NoAug and GreenAug-Rand, in one novel scene. The results, shown in Table 5 demonstrate that GreenAug-Rand applied to reinforcement learning with a different robot performs significantly better than NoAug.
Impact of texture randomness. We investigated how the texture randomness of GreenAug-Rand affects performance. We tested solid colours, Perlin noise (procedurally generated textures) [48], and MIL textures (used in the main experiments). All texture datasets are of the same size (5771). The evaluation was conducted on the “put cube in drawer” and “stack cups” tasks from the main experiment across three novel scenes (Scenes 3–5). Table 5 summarises the results. Consistent with domain randomization studies [25, 26, 27, 28, 29, 31], greater texture randomness leads to better performance. Examples of each texture type are provided in the Appendix.
@@ Appendix D Detailed Results
Appendix E Additional Limitations and Future Works

### v2

@@ 1 Introduction
In light of these complications, we opt for a simpler yet effective alternative: green screens. The film industry has utilised green screens extensively [14, 15, 16, 17, 18], enabling the addition of virtual backgrounds to live footage. Inspired by these applications, we apply green screen technology to robotics, allowing robots to perform tasks in unfamiliar scenes not part of the training data.
@@ 2 Related Work
Green screen in machine learning and robotics. Green screen has been traditionally used for film and video production [14, 15, 16, 17, 18]. In recent years, its application in machine learning has increased. Smirnov et al. [18] applied machine learning to improve the quality of chroma keying. Xu et al. [32], Sengupta et al. [33], Lin et al. [34, 35] explored machine learning techniques to replace green screens, enabling natural image matting without them. Schülein et al. [36] used green screens and chroma keying to replace backgrounds with clinical scenes to create synthetic data for medical clothing detection. In robotics, the use of green screens remains limited. Coates and Ng [37] used it to develop a multi-camera object detector with synthetic data from chroma-keyed backgrounds.
@@ 3.3 Training Robot Learning Policies
GreenAug can be applied to RGB-based robot learning methods. Similar to standard augmentation methods, images can be transformed with GreenAug and fed into policy networks during training, or they can be preprocessed offline and then used for training. Offline preprocessing is more common due to the longer computation time of some GreenAug variants. However, in online settings such as reinforcement learning, online transformations are also effective. GreenAug-Rand and GreenAug-Gen allow each raw frame from the training demonstrations to be augmented with different textures, significantly increasing the amount of preprocessed data. In contrast, GreenAug-Mask masks the background and only provides a one-to-one mapping of the original to augmented trajectory. To ensure a fair comparison, we keep the number of preprocessed frames equal to the number of raw frames for all methods.
@@ 4.2 Setup
Training. We trained all baselines and our methods on both sets of data, except for GreenAug, which was excluded from Scene 2 as it relies on the green screen. Each data set corresponds to a separate policy. ACT is used as the control policy for our main experiments. We use absolute joint positions as actions.
@@ 4.4 Ablation and Further Studies
Applying GreenAug to a different robot with reinforcement learning. We investigated whether GreenAug can be applied to a different robot embodiment and learning method, beyond the Franka Panda and ACT. We set up a similar “take lid off saucepan” task on a UR5. We used Coarse-to-fine Q-network [44] for the policy. The robot was provided with 24 demonstrations and was given a sparse reward of 0 for failure and 1 for success. We trained the robot online with 20 minutes of exploration on a green screen background and evaluated two policies, NoAug and GreenAug-Rand, in one novel scene. The results, shown in Table 5 demonstrate that GreenAug-Rand applied to reinforcement learning with a different robot performs significantly better than NoAug.
Impact of texture randomness. We investigated how the texture randomness of GreenAug-Rand affects performance. We tested solid colours, Perlin noise (procedurally generated textures) [45], and MIL textures [46] (used in the main experiments). All texture datasets are of the same size (5771). The evaluation was conducted on the “put cube in drawer” and “stack cups” tasks from the main experiment across three novel scenes (Scenes 3–5). Table 5 summarises the results. Consistent with domain randomization studies [25, 26, 27, 28, 29, 31], greater texture randomness leads to better performance. Examples of each texture type are provided in the Appendix.
@@ Appendix D Detailed Results
Appendix E Code Snippet for CVAug
The following code snippet outlines the data augmentation pipeline for CVAug. The pipeline begins by applying a random photometric distortion to the image. Next, it performs a random shift by cropping the image to its original size and adding a 4-pixel padding on each side. Each augmentation is wrapped in RandomApply, meaning each has a 50% chance of being applied. As a result, there’s a 50% chance of applying the photometric distortion, a 50% chance of performing the random shift, and a 25% chance of both augmentations being applied simultaneously.
import torchvision.transforms.v2 as TV
original_image_size = (240, 320)
TV.Compose(
TV.RandomApply([T.RandomPhotometricDistort()]),
TV.RandomApply(
[TV.RandomCrop(original_image_size, padding=4)]
), # random shift
Appendix F Additional Limitations and Future Works